OpenStudy (anonymous):

Given b = 4 and c = -5, evaluate b² - c².

OpenStudy (anonymous):

b^2-c^2 = (b+c) (b-c)

OpenStudy (anonymous):

b^2-c^2=(b+c)(b-c) I think you can do it now

OpenStudy (anonymous):

b^2-c^2 =(4)^2-(-5)^2 =16-25 =-9